OpenStudy (anonymous):

Two lines, C and D, are represented by the following equations: Line C: y = x + 5 Line D: y = −2x − 1 Which of the following options shows the solution to the system of equations and explains why? (−2, 3), because the point does not lie on any axis (−2, 3), because both lines pass through this point (−2, 3), because one of the lines passes through this point (−2, 3), because the point lies between the two axes

OpenStudy (cp9454):

B

OpenStudy (anonymous):

thanks so much do you mind helping me some more

OpenStudy (cp9454):

not at all

OpenStudy (anonymous):

Two lines, E and F, are represented by the equations given below. Line E: 5x + 5y = 40 Line F: x + y = 8 Which statement is true about the solution to the set of equations? It is (40, 8). It is (8, 40). There is no solution. There are infinitely many solutions.

OpenStudy (anonymous):

cp9454

OpenStudy (cp9454):

D, infinte many solution

OpenStudy (anonymous):

Line H is represented by the following equation: 2x + 2y = 8 What is most likely the equation for line K so the set of equations has infinitely many solutions? x + 2y = 2 x + y = 4 2x + y = 6 x − y = 8

OpenStudy (cp9454):

B

OpenStudy (anonymous):

\[\sum_{?}^{?}\]

OpenStudy (anguyennn):

Substitute y(0) for both equations and then x(0) for both of them as well, and it would give you a result for what the points may be.